OpenStudy (anonymous):

What is the y-intercept of line 2x plus 3y equals 6? Thanks.

OpenStudy (anonymous):

2x+3y=6 y= 2-(2/3)x The x-intercept is when y=0. Solving for y=0 (2/3)x=2 x=2(3/2) x=3 The slope is -(2/3) (y2-y1)/(x2-x1) x=0 y=2, y=0 x=3 (0-2)/(3-0) = -2/3;)